### v2

# The Overcooked Generalisation Challenge

Abstract
We introduce the Overcooked Generalisation Challenge (OGC) – the first benchmark to study reinforcement learning agents’ zero-shot cooperation abilities when faced with novel partners and levels in the Overcooked-AI environment. This perspective starkly contrasts a large body of previous work that evaluated cooperating agents only on the same level or with the same partner, thus failing to capture generalisation abilities essential for real-world human-AI cooperation. Our challenge interfaces with state-of-the-art dual curriculum design (DCD) methods to generate auto-curricula for training general agents in Overcooked. OGC is the first environment specially designed for DCD methods and, consequently, the first evaluated with state-of-the-art methods. It is fully GPU-accelerated, built on the DCD benchmark suite minimax, and freely available as open source. We show that state-of-the-art DCD algorithms fail to produce useful policies on OGC, even if combined with recent network architectures specifically designed for scalability and generalisability. As such, the OGC pushes the boundaries of real-world human-AI cooperation by enabling pioneering research on the impact of generalisation on cooperating agents.
Human-AI Cooperation, Multi-Agent Reinforcement Learning, Unsupervised Environment Design, Benchmark
1 Introduction
Figure 1: The Overcooked Generalisation Challenge (OGC) is the first benchmark that requires reinforcement learning agents to cooperate with novel partners in novel levels in a zero-shot fashion. During training, agents can access a generator that outputs new training environments. During evaluation, agents are presented with a novel level and partner to cooperate with.
Developing computational agents capable of collaborating with humans has emerged as a key challenge in artificial intelligence (AI) research (Stone et al., 2010; Dafoe et al., 2020) and promises to vastly expand human abilities (O’neill et al., 2020). Recent years have seen considerable advances in understanding human cooperative behaviour (Rand & Nowak, 2013; Vizmathy et al., 2024), computational modelling of cooperation (Nikolaidis & Shah, 2013; Sadigh et al., 2016; Ding et al., 2024), as well as in developing computational methods for human-AI cooperation (Hu et al., 2020; Strouse et al., 2021). In parallel, several benchmarks were proposed to foster the development and evaluation of these methods (Samvelyan et al., 2019; Bard et al., 2020). Most notably, Overcooked-AI (Carroll et al., 2019) has established itself as a widely used benchmark for evaluating (zero-shot) human-AI coordination (Strouse et al., 2021; Zhao et al., 2023; Yu et al., 2023).
Despite the advances they have enabled, all of these benchmarks are limited in that they only allow to assess reinforcement learning (RL) agents’ cooperative abilities in-distribution. That is, they either only allow to evaluate agents in the same environment in which they were trained (Hu et al., 2020; Carroll et al., 2019) or the same partner agent they were trained with (Foerster et al., 2018; Lowe et al., 2017; Strouse et al., 2021). In Overcooked-AI, for instance, existing zero-shot coordination (ZSC) methods are trained once per layout at considerable computational cost (Carroll et al., 2019; Yang et al., 2022; Zhao et al., 2023; Yu et al., 2023), and these layouts only feature a limited number of possible cooperation strategies. However, real collaborative settings require coordination with novel partners and in unknown environments. While previous open-source benchmarks studied opponents combined with map generalisation, no dedicated open-source benchmarks exist for studying cooperation partners combined with map generalisation. Consider a medical robot assisting doctors in hospitals as an example of why this might be useful. Such a robot will be deployed in unique and unknown hospitals and surgical rooms, where it must adapt to different medical staff and their preferences.
To address this limitation, we introduce the Overcooked Generalisation Challenge (OGC) – the first zero-shot cooperation benchmark that challenges agents to cooperate in novel layouts and with unknown partners, see Figure 1. Cooperative settings differ from competitive ones in their game-theoretic background and thus require separate algorithms and benchmarks (Lerer & Peysakhovich, 2019). Neural MMO (Suarez et al., 2021, 2023) comes closest as it mixes cooperation and competition but – unlike Overcooked-AI – does not provide a purely cooperative setting that is specifically designed for human-AI collaboration.
To train and evaluate agents on OGC, we make use of unsupervised environment design (UED) (Dennis et al., 2020) to generate suitable training levels, provide hand-designed testing levels, assess ad-hoc teamwork on these by providing populations of diverse testing agents and then similarly also test for zero-shot cooperation. As such, our work is the first to combine UED techniques with a multi-agent RL zero-shot cooperation task and thus bridges the gap between two previously unrelated research areas; it studies the impact of generalisation on human-AI coordination and the ability of UED algorithms to design optimal auto-curricula for cooperating agents. We show that several state-of-the-art UED algorithms and network architectures struggle to perform well on our challenge. Only PAIRED (Dennis et al., 2020) together with a policy that incorporates a soft Mixture-of-Experts (SoftMoE) module (Obando-Ceron et al., 2024) has limited success at generalising to the testing levels and outperforms competitive baselines, including robust PLR (Jiang et al., 2021b, a) and ACCEL (Parker-Holder et al., 2022). Our findings call for future work to develop new methods combining zero-shot coordination and DCD techniques in a single ZSC-DCD framework. Our benchmark provides the necessary environment to facilitate these developments. Taken together, our contributions are three-fold:
  1. 1.
    We introduce the Overcooked Generalisation Challenge – a novel benchmark challenge in which agents are asked to cooperate with novel partners and in previously unseen layouts.
  2. 2.
    We provide OvercookedUED – an open-source environment that can be used with state-of-the-art DCD algorithms and that is integrated into minimax (Jiang et al., 2023), taking full advantage of hardware acceleration provided by JAX.
  3. 3.
    We establish baseline performance on the OGC using state-of-the-art DCD algorithms and neural network architectures (Smith et al., 2023; Obando-Ceron et al., 2024). We show that current DCD algorithms struggle with our new challenge, including the most recent network architectures. Notably, we show the generalisation performance of SoftMoE-based architectures in hard generalisation tasks. Furthermore, we assess zero-shot cooperation performance with a population of diverse partners to link zero-shot cooperation and generalisation. We show that, as policies become more generally capable, they achieve better zero-shot cooperation.
2 Research Challenges
Figure 2: The OGC features a challenging large design space where many different elements have to be placed in relation to each other. This makes the OGC challenging for both designing learning levels and training agents.
The OGC poses several new challenges for zero-shot human-AI cooperation that go beyond existing benchmarks and that are essential for further advances in the development of cooperating RL agents:
Generalisation
The OGC challenges the generalisation capabilities of methods and agents by having them engage in a double generalisation challenge: adjusting to both novel partners and levels. Existing cooperative open-source benchmarks typically require only one form of generalisation; see, for instance, (Lowe et al., 2017; Foerster et al., 2018; Carroll et al., 2019; Hu et al., 2020).
Environment Design
Our environment challenges UED algorithms in generating and designing layouts with many interacting components and agents. This is in contrast to existing environments that only require UED algorithms to design simple mazes, 2D walker terrains, or race tracks consisting of fewer elements (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022; Rutherford et al., 2024a). We show that methods struggle to design layouts similar to the ones humans designed. Current methods specifically fail to design layouts requiring handing over items over a countertop or featuring deliberately designed circuits. Our benchmark challenges further research to develop UED methods that design more realistic collaboration environments for curriculum learning, possibly along the lines of Bruce et al. (2024).
Combining Environment and Partner Generalisation
Coordinating with novel partners and generalising to novel levels were often treated as separate tasks. Previous work on methods for zero-shot coordination often requires training a diverse population of partners. Since there currently is no algorithm to train a population of diverse agents over a distribution of levels, these methods need adjusting to be applicable to our challenge. Our benchmark thus encourages merging both lines of research and developing UED-ZSC methods, i.e. methods that learn both at the same time, in a single joint generalisation task.
3 Related work
3.1 Generalisation in Reinforcement Learning
A large number of works have shown that RL agents fail to generalise to new environments (Zhang et al., 2018a; Cobbe et al., 2019) and have triggered research on the generalisation capabilities of RL agents (Nichol et al., 2018; Cobbe et al., 2019, 2020). Early results revealed that RL agents can memorise large numbers of levels during training (Zhang et al., 2018b; Cobbe et al., 2019) and that they must experience sufficiently diverse training data to generalise well (Cobbe et al., 2020). One established approach to generate diverse training data is to use domain randomisation (Jakobi, 1997, DR). Still, DR has been shown to produce many uninformative samples (Khirodkar et al., 2018), which can lead to the agent’s inability to generalise (Dennis et al., 2020).
3.2 Unsupervised Environment Design
Intending to address this challenge, later works on generalisation focused on unsupervised environment design (Dennis et al., 2020, UED). UED aims to improve domain randomisation by generating auto-curricula that include training levels of increasing complexity to facilitate continued agent learning (Graves et al., 2017). It does so by adapting the free parameters of an under-specified environment to the agent’s capabilities. Most popular UED methods fall into the category of Dual Curriculum Design (Jiang et al., 2021a, DCD) that combine 1) an agent, 2) a level generator, and 3) a curator that picks which levels to train on. Popular methods include Prioritised Level Replay (PLR) (Jiang et al., 2021b), robust PLR (Jiang et al., 2021a), MAESTRO (Samvelyan et al., 2023), ReMiDi (Beukman et al., 2024), PAIRED (Dennis et al., 2020), ACCEL (Parker-Holder et al., 2022), and Replay-Enhanced (RE)PAIRED (Jiang et al., 2021a). While the development of these DCD methods has been steady, they have mostly been explored in simple single-agent environments For a comprehensive overview, see Table 3 and Appendix A.3. Compared to single-agent environments, multi-agent environments are inherently more complex because the agents interact with each other, as well as with the environment. Multi-agent environments are still rarely used in UED research. Most prominent is Deepmind’s XLand (Team et al., 2021; Bauer et al., 2023), a closed-source multi-task universe for generating single- and multi-agent tasks and environments. While XLand features cooperative tasks, it is not available to researchers for studying cooperative multi-agent UED algorithms. While an open-source variant was recently published (Nikulin et al., 2023), it only supports a single agent. Arguably closest is Neural MMO (Suarez et al., 2021, 2023), a massively multi-task and multi-agent environment that combines cooperation and competition to replicate massively multiplayer online games. Instead, we are interested in assessing and identifying cooperation performance in specially designed human-AI cooperation challenges for which the massive multi-tasking and multi-agent cooperation-competition setting of Neural MMO is unsuitable. Additionally, classic Overcooked benefits from a rich history of human-AI cooperation research. Finally, while JaxNav (Rutherford et al., 2024a) features multi-agent path-finding, no interaction between agents is required, and the environment is not focused on human-AI cooperation. Other open-source environments are competitive, i.e. LaserTag (Lanctot et al., 2017; Samvelyan et al., 2023) and MultiCarRacing (Schwarting et al., 2021; Samvelyan et al., 2023), and thus not applicable to our setting. Opposed to all of these, our work contributes to and analyses the first open-source cooperative multi-agent UED environment.
3.3 Human-AI Cooperation in Overcooked-AI
Overcooked-AI (Carroll et al., 2019) has become one of the most important benchmarks for human-AI cooperation. The environment is fully cooperative and has two agents cook and deliver soups to earn a joint reward. Overcooked-AI was, for example, used in research on zero-shot cooperation (Strouse et al., 2021; Zhao et al., 2023; Yu et al., 2023; Li et al., 2023b; Yan et al., 2023, ZSC), language model-based cooperative agents (Liu et al., 2024; Tan et al., 2024), human modelling in cooperation (Yang et al., 2022). Zero-shot cooperation (Hu et al., 2020) refers to cooperating with a partner not encountered during training. It is an important proxy to ensure the ability of an agent to coordinate with humans at test time, given that human data is often unavailable and agents thus must be able to coordinate effectively without previous training. It is commonly studied in Overcooked. Our work is thus the first to explore the impact of cross-level generalisation for zero-shot cooperation and is the first to provide the necessary tools for this.
4 Preliminaries
The cooperative multi-agent UED setting can be formalised as a decentralised under-specified partially observable Markov decision process (Dec-UPOMDP) with shared rewards. A Dec-UPOMDP is defined as $\mathcal{M}=\langle\mathcal{N},A,\Omega,\Theta,\mathcal{S}^{\mathcal{M}},\mathcal{T}^{\mathcal{M}},O^{\mathcal{M}},\mathcal{R}^{\mathcal{M}},\gamma\rangle$ in which $\mathcal{N}$ is the set of agents with cardinality $n$, $\Omega$ is a set of observations, and $\mathcal{S}^{\mathcal{M}}$ is the set of true states in the environment. Partial observations $o^{i}\in\Omega$ are obtained by agent $i\in\mathcal{N}$ using the observation function $O:\mathcal{S}\times\mathcal{N}\rightarrow\Omega$. Following (Jiang et al., 2021a), a level $\mathcal{M}_{\theta}$ is defined as a fully-specified environment given some parameters $\theta\in\Theta$. In it, agents each pick an action $a_{i}\in A$ simultaneously to produce a joint action $\boldsymbol{a}=(a_{1},\dots,a_{n})$ and observe a shared immediate reward $R(s,\boldsymbol{a})$. Then, the environment transitions to the next state according to a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}^{1}\times...\times\mathcal{A}^{n}\times\Theta\rightarrow\Delta(\mathcal{S})$ where $\Delta(\mathcal{S})$ refers to the space of distributions over $\mathcal{S}$. $\gamma\in[0,1)$ specifies the discount factor. Agents learn a policy $\pi$. The joint policy $\boldsymbol{\pi}$ together with the discounted return $R_{t}=\sum^{\infty}_{i=0}\gamma^{i}r_{t+1}$ induce a joint action value function $Q^{\boldsymbol{\pi}}=\mathbb{E}_{s_{t+1:\infty},\boldsymbol{a}_{t+1:\infty}}[R_{t}|s_{t},\boldsymbol{a}_{t}]$. The definition of the Dec-UPOMDP extends a Dec-POMDP (Oliehoek & Amato, 2016; Wu et al., 2021) with the free parameters of the environment $\Theta$, analogously to previous works (Dennis et al., 2020; Jiang et al., 2021a; Samvelyan et al., 2023).
Our definition differs from (Samvelyan et al., 2023) in terms of the shared rewards and general-sum nature. Within our Dec-UPOMDP, we perform UED to train a policy over a distribution of fully specified environments that enable optimal learning. This is facilitated by obtaining an environment policy $\Lambda$ (Dennis et al., 2020) that specifies a sequence of environment parameters $\Theta^{T}$ for the given policy that is to be trained. How $\Lambda$ is obtained depends on the DCD method. In OvercookedUED, $\Theta$ represents the possible positions of walls, pots, serving spots, agent starting locations, and onion and bowl piles adjusted by $\Lambda$ throughout training.
5 The Overcooked Generalisation Challenge
Figure 3: Overview of the OGC and how it is typically used in a DCD algorithm. The OGC supports teacher-based UED methods like PAIRED (Dennis et al., 2020) and edit-based methods like ACCEL (Parker-Holder et al., 2022) via mutator functions of existing layouts.
Figure 3 provides an overview of the Overcooked Generalisation Challenge. The OGC extends previous work by evaluating the cooperative abilities out-of-distribution. In contrast to existing UED environments, the OGC focuses on the cooperation of multiple agents in a complex, cooperative task across different levels. More specifically, two different agents are tasked with cooking a soup together in the five original layouts of Overcooked-AI (see Figure 1 bottom) but without encountering them or their partner during training. The original five layouts have been designed to test and explore different kinds of cooperation and thus form suitable out-of-distribution test levels. Since the OGC does not limit evaluation to one specific physical environment or partner, it is more closely aligned with real-world human-AI collaboration. To train an agent capable of generalisation, we generate a curriculum of training layouts via procedural content generation. To generate a curriculum of layouts, we use DCD methods. Specifically, methods in which an environment designer interacts with the challenge by designing layouts either from scratch through interacting with OvercookedUED – a novel environment for creating Overcooked levels – by alternating existing layouts through the Overcooked mutator or by letting the OGC generate random layouts. At every step of the curriculum, this designer must account for agents’ cooperation ability when trying to generate layouts that are at the forefront of their abilities. While OGC refers to the challenge as a whole, it comprises several components that enable its integration with DCD algorithms (see Figure 3). At the heart, it features an Overcooked environment capable of running different levels fast and in parallel in which agents learn to cooperate. It features OvercookedUED that provide methods, interfaces and a teacher environment to design novel layouts and an Overcooked Mutator that alters existing layouts.
Overcooked-AI
OGC builds on the Overcooked-AI environment provided by the JaxMARL project (Rutherford et al., 2024b), keeping features consistent with the original implementation. This includes action and observation spaces, i.e. the set of actions is {left, right, up, down, interact, stay} and observations are encoded as a stack of 26 $h\times w$ boolean masks encoding the positions of elements in the environment. In this representation, the first mask encodes the position of the first agent; the second mask is the one of the second agent, etc. We extended their Overcooked implementation to be parallelisable over different layouts – which is impossible in the JAXMarl version – to train agents on many different layouts simultaneously and efficiently. This requires padding all layouts to the same size. I.e., during rollouts, layouts are padded to a maximum size, and all objects in these layouts are one-hot encoded based on their position in equally sized masks. While this facilitates fast parallel rollouts that can be compiled just-in-time, it requires the introduction of a maximum height $h$ and width $w$ that need to be picked as a hyper-parameter before training.
OvercookedUED
OvercookedUED features the interfaces necessary to design new layouts. For algorithms that make use of a teacher agent to create layouts (PAIRED, etc.), OvercookedUED provides a teacher environment (see Figure 3 right). This environment allows a teacher policy to take design steps to parametrise the underspecified MDP. At every timestep $t$ of the generation process, the teacher observes the unfinished layout and picks an action from a space that consists of the total number of cells in the $h\times w$ grid. This cell then becomes filled with the next item to be placed. Objects are placed sequentially and in a deterministic order, starting from walls, agents one and two, goal, onion, pot and bowl positions. An episode in the teacher MDP lasts until all items are placed. When placing a new element on top of an existing one of a different type, the new element is placed randomly on a free cell. Placing two elements of the same type on the same cell results in the second not being placed. This allows designers to place various elements per type, e.g. one to two onions. This is consistent with prior work (Dennis et al., 2020). As in previous work (Jiang et al., 2023), OvercookedUED does not check whether a layout is solvable and leaves the task of designing and/or identifying suitable training layouts to the DCD method. For algorithms that do not specify a teacher, such as PLR, OvercookedUED generates random layouts. These random layouts feature one or two piles of onions, bowls, pots and serving locations, and two agents. Finally, some DCD algorithms, such as ACCEL, require alternating existing layouts by mutating them. OvercookedUED supports layout mutation through five basic operations: (1) converting a random wall to a free space and vice versa, (2) moving goals, (3) pots, (4) plate piles, and (5) onion piles. Our mutator does not move agents. Given a layout, our mutator randomly samples $n$ operations and applies them. All versions allow the number of walls to be configured, and the environment always places a border wall.
Figure 4: The SoftMoE-LSTM agents architecture used in this work. We employ the PerConv tokenisation technique introduced in (Obando-Ceron et al., 2024).
Evaluation
We evaluate agents by their performance on out-of-distribution Overcooked-AI layouts to asses generalisation performance in self-play, ad-hoc teamplay (Stone et al., 2010) and zero-shot coordination. In ad-hoc teamplay, a fictitious co-play (Strouse et al., 2021, FCP) and maximum entropy-based population-based training (Zhao et al., 2023, MEP) population of a total 24 agents each is used to asses cooperation performance with novel partners. Both populations are trained with equal settings and include a low, medium and high-skilled checkpoint of each run extracted at 10%, 50%, and 100% achieved return, respectively. The population entropy coefficient $\alpha$ is $0.01$ for MEP. We use cross-play with a diverse population and zero-shot cooperation as a proxy for assessing the abilities of an agent to coordinate with humans. In zero-shot cooperation, we task agents to coordinate with different random seeds of the same method (Hu et al., 2020). Our benchmark includes the necessary tools to perform this evaluation. In our analysis, we report results using the mean episode reward and mean layout solved rate, similar to previous work (Jiang et al., 2023). A layout is considered solved if an agent pair delivers more than one soup, differentiating goal-directed from random behaviour. We present these metrics in the self- and cross-play settings. Additionally, we investigate what kinds of levels agents perform poorly in and why in a final error analysis.
6 Analysing & Benchmarking
Figure 5: Return of policies in all evaluation layouts. Left: Results of policies trained across all five evaluation layouts to be used as an oracle. SoftMoE-LSTM shows the best performance. Right: Shows SoftMoE-LSTM agents trained only on one layout, which overfit.
Figure 6: Sample levels generated by the different methods after $15,000$ (Middle) and $30,000$ (End) epochs. Even after considerable training, none of the methods can guarantee the generation of solvable layouts (Middle-row leftmost and rightmost).
We establish baseline performance on the OGC using state-of-the-art architectures, DCD algorithms, and neural network architectures. We aim to set a performance baseline for future works and show what evaluations this benchmark enables. To this end, we first show that generalising to novel layouts in Overcooked is difficult, and then we move on to the additional challenge of zero-shot cooperation.
All baselines are trained using MAPPO, which works well in cooperative settings (Yu et al., 2022). As for DCD algorithms, we compare the performance of DR, PLR⟂,∥, Pop. PAIRED and ACCEL. We chose these methods as they have better theoretical guarantees (PLR vs PLR), better runtime performance (ACCEL and PLR), or because we found them to perform better empirically (Pop. PAIRED vs PAIRED). We excluded POET (Wang et al., 2019) in this analysis as it outputs specialists rather than generalists, which we require (Parker-Holder et al., 2022). Additionally, we excluded MAESTRO as it is based on prioritised fictitious self-play (Heinrich et al., 2015; Vinyals et al., 2019) that is not readily adaptable to the cooperative setting (Strouse et al., 2021). In our experiments, the number of walls to place for a specific layout is initially drawn randomly (PLR, ACCEL, DR) or chosen by the action of the teacher agent (PAIRED) from the interval $[1,15]$ where $15$ is the configured maximum. From there on, walls are placed at random (PLR, ACCEL, DR) or based on the actions of the teacher agent (PAIRED), which is parametrised by its own neural network. All methods place at least one and up to two items per category. Since ACCEL can add walls through editing, we restrict ACCEL from adding more walls than the maximum configured to keep methods comparable. In this case, ACCEL can only remove walls or move items. The same holds for PAIRED, where we pick the number of design steps so that no more than the maximum number of walls can be set. The design space of all methods is thus the same.
As in (Jiang et al., 2023), if not stated otherwise, we train in $32$ parallel environments and stop after $30,000$ outer training loops, amounting to just under $400$ million steps in the environment. Hyperparameters were picked after a grid search over reasonable values, and all parameters are provided in Appendix A.4.
Neural Architectures and Agents
Table 1: Mean episode reward for the different methods averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over three random seeds. We include Oracles trained on the five testing layouts to establish an empirical upper bound.
Method CNN-LSTM SoftMoE-LSTM CNN-S5
DR $0.46\pm 0.16$ $5.22\pm 7.19$ $00.00\pm 0$ 0.00
PLR $0.17\pm 0.06$ $0.91\pm 0.71$ $0.12\pm 0.15$
PAIRED $0.64\pm 0.38$ $\mathbf{9.52\pm 1.02}$ $00.00\pm 0$ 0.00
ACCEL $0.40\pm 0.35$ $0.72\pm 0.59$ $0.09\pm 0.12$
Oracle $\mathbf{189.49}\pm 12.96$ $\mathbf{217.02}\pm 39.18$ $\mathbf{155.01}\pm 12.82$
Figure 7: Returns over training for both training and evaluation layouts for our SoftMoE-LSTM-PAIRED policy. The policy has had some success in generalising, but its generalisation gap remains substantial.
We first compared suitable network architectures for our agents. Our default neural network architecture consists of a convolutional encoder with a recurrent neural network with an LSTM (Hochreiter & Schmidhuber, 1997). We picked it because of its performance in previous work (Yu et al., 2023). In addition, we also explore the use of SoftMoE (Obando-Ceron et al., 2024), which have recently been identified for their potential for enabling scaling and generalisation, and S5 (Smith et al., 2023) layers due to the strong results of structured state-space models (Gu et al., 2022) in meta reinforcement learning (Lu et al., 2023). SoftMoE modules replace the penultimate layer after the feature extractor and S5 layers the LSTM in all experiments. For SoftMoE we use the PerConv tokenisation technique (Obando-Ceron et al., 2024). We hypothesise that these provide better generalisation performance. All details are in Appendix A.5.
We first compare their ability to fit all evaluation layouts simultaneously. We refer to these agents as Oracles and use them later for comparison. As presented in Figure 5 (left), all agents can fit all five layouts simultaneously. Note that the S5 agent underperforms considerably while the SoftMoE-LSTM agent performs best, which is why we will use them in future experiments. As a baseline, we compare against a SoftMoE-LSTM agent that is only fitted to one layout and evaluated on all (Figure 5 (right)) which overfits. Because of our finetuned SoftMoE model and implementation, we see very high self-play returns compared to the literature. The SoftMoE-LSTM Oracle sometimes even outperform the overfitting baseline. Detailed results are in Figure 5 and training curves are in Appendix A.7.1.
Layout Generalisation Performance
Despite our fine-tuned implementation, simply generalising to the testing layouts in the OGC is already challenging for all methods without coordinating with novel partners, as presented in Table 1. Compared to commonly used single-agent Maze environments such as AMaze, all DCD methods struggle to obtain good results. This is most evident when compared with oracle policies (bottom row). Since the environment features many moving parts to be placed correctly (see Figure 2), approaches like DR have difficulties in finding optimal placements by chance, as reflected in the results. PAIRED outperforms all other models significantly $p<0.05$ using a one-sided paired t-test. This is also shown in the mean solved rate where it reaches $14.6\pm 7.7\%$, while all other models have a solved rate of mostly $0\%$ (cf. Appendix A.6.2). While this model performs better on average, layouts differ greatly in their difficulty. Our best-performing model reaches modest performance in Asymmetric Advantages and Cramped Room while mostly failing in others, with no other model achieving noteworthy results. We show the training and evaluation return of SoftMoE-LSTM-PAIRED in Figure 7 and 15. The curve shows the model converging but never generalising to the test levels. The full results are in the Appendix A.6.3.
Performance with Novel Partners
Figure 8: Ad-hoc team-play performance of SoftMoE-LSTM-PAIRED and several baselines with both an MEP and an FCP expert population. While SoftMoE-LSTM-PAIRED outperforms simple baselines, it fails to compete against the oracle.
Figure 9: Zero-shot cooperation performance of both the oracle and the SoftMoE-LSTM-PAIRED agents (trained with different seeds).
Ultimately, we want the OGC to connect map and partner generalisation and thus evaluate both ad-hoc team play and zero-shot coordination. For ad-hoc team play, we train and then use an FCP and MEP population to establish how our and future agents can coordinate with diverse policies. See Appendix A.6.4 for details. We present results for the SoftMoE-LSTM-PAIRED policy in Figure 8 together with three other baselines: stay, which is a partner that never moves, random, which samples random actions and oracle as discussed above. Results for all methods are in the Appendix in Figure 13. As performance on out-of-distribution levels rises, agents become more competent at coordinating. PAIRED always outperforms baselines (cf. Appendix A.6.5). However, even PAIRED policies often perform only slightly better than random baselines, which signifies the challenges of our benchmark. This is also evidenced by the kinds of levels these methods generate (Figure 6), as they tend to pivot towards generating open spaces that ease cooperation but are notably different from evaluation layouts. Overall, cooperation performance is carried mainly by the expert FCP and MEP agents (compare Table 13 and 14), mostly since our agents struggle to perform on the evaluation layouts in the first place.
Additionally, to evaluate our agents’ full spectrum of out-of-distribution abilities, we establish zero-shot cooperation performance with agents themselves over three random seeds. We display results in Figure 9. Two results stand out in the Figure. First, our SoftMoE-LSTM-PAIRED agents perform badly in both self- and cross-play in the three harder evaluation layouts (Coordination Ring, Forced Coordination and Counter Circuit). Second, in the two remaining easier layouts, our agent sometimes performs better with a novel partner than with itself. The same is never true for the oracle baseline. While this does not present definite evidence, it hints at a connection between map generalisation and zero-shot cooperation performance in Overcooked.
Error Analysis
Figure 10: Sample levels that our models perform best (top) and worst in (bottom). The number of visits to each grid cell is shown as a heatmap overlay, while the mean return is below. The layouts where the model performs worst tend to feature narrow elements or large distances between items.
We perform two final experiments to investigate our baselines’ poor performance and eliminate trivial sources of errors. First, we hypothesise that the top-down observations in OGC make generalisation difficult as they are not invariant to mirroring or rotations (Ye et al., 2020). To test this, we evaluated agents on 24 hand-designed circular evaluation levels with different kinds of symmetry, as shown in Figure 14. We found that agents tend to perform similarly across these layouts as the standard deviation is at most $1.1$, and therefore reject this hypothesis (more details in Appendix A.7). Second, we investigate the kinds of levels our best-performing model does well vs poorly in from a pool of randomly generated evaluation levels in Figure 10. The figure summarises the cooperation behaviour of the agents by showing which cells are visited most frequently to give an impression of their motion patterns. While on many layouts, our model reaches good self-play performance (up to a maximum mean reward of $84.4$; top row), it typically delivers few to no soups in layouts that it performs worst in. These levels tend to be narrow/convoluted and/or feature big distances between objects. In conclusion, current DCD methods struggle with generating training layouts of correct complexity, i.e. ones that are similarly hard to evaluation ones.
Discussion
Previous work (Jiang et al., 2021a) found that PLR outperforms the other here-tested algorithms in navigation-based tasks. Our more challenging environment suggests that this might not always be the case. In our preliminary analysis, PAIRED outperformed other DCD methods. Compared to mazes, car racing, or walker environments with fewer moving pieces, Overcooked layouts are more complex to design, requiring the designer to place multiple objects in relation to each other and the agents (see Figure 2). Therefore, methods that employ a random generator struggle in such a big design space. This thus requires a capable generator and suggests that simple navigation-based environments used to benchmark DCD in UED algorithms do not allow full performance evaluation. As such, OvercookedUED can be an important part of evaluating DCD algorithms. We envision that general Overcooked agents should be evaluated in scenarios that are difficult for self-play agents using our benchmark. These include coordinating with strongly-biased agents (Yu et al., 2023) in Coordination Ring (see Section 1) and Asymmetric Advantages as described in (Ruhdorfer, 2023) and for which we provide the tools.
Table 2: Performance on mirrored and rotated levels, illustrated in Figure 14. The low variance suggests that the agents do not struggle with spatial reasoning.
Method SoftMoE-LSTM CNN-S5 CNN-LSTM
DR $11.91\pm 0.8$ $0.00\pm 0$ 0.0 $1.96\pm 0.3$
PLR $4.39\pm 0.4$ $0.49\pm 0.2$ $0.65\pm 0.2$
PAIRED $36.83\pm 1.1$ $0.58\pm 0.2$ $0.59\pm 0.1$
ACCEL $2.36\pm 0.3$ $1.69\pm 0.2$ $0.78\pm 0.2$
7 Limitations
Despite its many advantages, our challenge has two main limitations. First, we artificially restricted the maximum size of the layouts to allow the environment. Future work should focus on more natural representations of the whole scene, e.g. using graphs or item embeddings. While we included a partial observation that could theoretically be computed independently of size, similar to the vector-based observation used for behaviour cloning agents in (Carroll et al., 2019), batching across layouts in OvercookedUED still requires the layouts to be scaled to the same height and width. Second, while our challenge allows us to study zero-shot coordination via generalising across layouts, reasoning about other agents (Rabinowitz et al., 2018; Gandhi et al., 2021; Bara et al., 2023; Bortoletto et al., 2024b, a) might be equally important to achieve zero-shot cooperation capabilities on unknown layouts. This is plausible given that humans can reason about the mental states of other agents via Theory of Mind (Premack & Woodruff, 1978), as well as the physical configuration of the space in which they operate. Future work could thus explore reasoning about other agents in previously unexplored environments.
8 Conclusion
We presented the Overcooked Generalisation Challenge (OGC) – a generalisation challenge focusing on (zero-shot) cooperation in MARL in out-of-distribution test levels. Our challenge is the first open-source cooperative multi-agent UED environment, which is significantly more challenging than previous environments commonly used in UED and DCD research. In addition to using the challenge in UED research, we have shown how the OGC can be used in future research on human-AI collaboration as a zero-shot cooperation benchmark for general agents. That is, our challenge establishes a link between generalisation and zero-shot coordination. Our work is the first to provide the research community with the tools to train and evaluate agents capable of coordinating in previously unknown physical spaces and with novel partners.
Impact Statement
While our work is largely foundational and concerned with providing the research community with the appropriate tools for the training and evaluation of agents in game-like environments, special caution is always imperative should this research be applied to human-AI collaboration. Even though our goal is to improve collaboration, safeguards should be applied to make sure that humans are always safe from harm. Especially in real-world applications where accidents could potentially result in bodily harm. Since our work is still far removed from any real-world application, we do not expect that our work in its present form carries the risk of materialising these harms. Some form of unsupervised environment design in collaborative environments might be part of future systems, and we, therefore, acknowledge these risks. This work, of course, also carries the potential to improve human-AI collaboration, and we make an important contribution to advancing the field with potential impacts in all kinds of human-machine interaction.
Appendix A Appendix
A.1 Accessibility of the benchmark
We make our challenge available under the Apache License 2.0 via a code repository: https://git.hcics.simtech.uni-stuttgart.de/public-projects/OGC. Our environment is built on top of the existing minimax project (accessible under Apache License 2.0 via https://github.com/facebookresearch/minimax) and is thus accessible to researchers who are already familiar with the project. minimax is extensively documented, fast, and supports multi-device training. For all details, including a full description of the advantages of minimax, we kindly refer the reader to the accompanying publication (Jiang et al., 2023). Our Overcooked adaption is extended from the one in JaxMARL also accessible under Apache License 2.0 via https://github.com/FLAIROx/JaxMARL. Our code includes extensive documentation and examples of how it may be used. Additionally, our code is written in a modular fashion and other multi-agent environments can be integrated with the runners.
A.2 Infrastructure & tools
We ran our experiments on a server running Ubuntu 22.04, equipped with NVIDIA Tesla V100-SXM2 GPUs with 32GB of memory and Intel Xeon Platinum 8260 CPUs. All training runs are executed on a single GPU only. We trained our models using Jax (Bradbury et al., 2018) and Flax (Heek et al., 2023) with 1, 2 and 3 as random seed for training DCD methods and 1 to 8 as random seeds for the populations. Training the DCD methods usually finishes in under 24 hours, only SoftMoE and PAIRED-based methods take longer. SoftMoE-based policies often take an extra 50% wall-clock time to train. Noticeable is also that our S5 implementation is the fastest, usually needing roughly 30% less time. Both are compared to the default architectures’ training time. In the longest case, the combination of a SoftMoE-LSTM policy trained with PAIRED takes about 80 hours to complete training. Our benchmark should be runnable on any system that features a single CUDA-compatible GPU. Although in our experience our experiments will require 32GB VRAM to run.
A.3 Extended Related Work
Table 3: Overview of benchmarks for unsupervised environment design and procedurally generated environments. Closed-source benchmarks are marked in gray – these cannot be evaluated on by the research community.
Name Multi-agent Zero-shot coop. GPU accelerated Open Source Partial obs. Img. obs.
XLand (Team et al., 2021; Bauer et al., 2023) - ?
LaserTag (Samvelyan et al., 2023) - - -
MultiCarRacing (Samvelyan et al., 2023) - - -
CoinRun (Cobbe et al., 2019) - - -
ProcGen (Cobbe et al., 2020) - - -
2D Mazes (Cobbe et al., 2019; Dennis et al., 2020) - - -
CarRacing (Jiang et al., 2021a) - - -
Bipedal Walker (Wang et al., 2019) - - - -
AMaze (Jiang et al., 2023) - -
XLand-MiniGrid (Nikulin et al., 2023) - -
Craftax (Matthews et al., 2024) - -
JaxNav (Rutherford et al., 2024a) - -
OvercookedUED (ours)
We present an overview over how our environment compares to other UED environments in Table 3. Additionally, we compare our work to single-agent UED environments and other attempts at using procedural contentment generation for Overcooked-AI below.
Single-Agent UED Environments
Early work on generalisation mainly focused on single-agent environments (Zhang et al., 2018b; Farebrother et al., 2018; Cobbe et al., 2019) and these are also popular in UED research. Among these, prior work has studied mazes (Dennis et al., 2020; Jiang et al., 2021a; Parker-Holder et al., 2022; Jiang et al., 2023; Li et al., 2023a; Beukman et al., 2024), bipedal walkers (Wang et al., 2019, 2020; Parker-Holder et al., 2022) or car racing environments (Jiang et al., 2021a). One likely reason for their popularity as benchmarks for DCD is that new levels are easy to generate, and agents are usually fast to train. However, they are limited to a single agent, with limited options to interact with the environment and other agents, and thus bear little resemblance to real-world problems.
Procedual-Content Generation in Overcooked
Related to our work is the work of (Fontaine et al., 2021) in which the authors used procedurally generated Overcooked layouts to evaluate the impact of different layouts on human-robot interaction using planning algorithms. However, while they use procedural content generation in the Overcooked context their research does not focus on cross-layout generalisation – a major theme in our work.
A.4 Hyperparameters
Table 4: Hyperparamters of the learning process.
Description Value
Optimizer Adam (Kingma & Ba, 2015)
Adam $\beta_{1}$ $0.9$
Adam $\beta_{2}$ $0.999$
Adam $\epsilon$ $1\cdot 10^{-5}$
Learning Rate $\eta$ $3\cdot 10^{-4}$
Learning Rate Annealing -
Max Grad Norm $0.5$
Discount Rate $\gamma$ $0.999$
GAE $\lambda$ $0.98$
Entropy Coefficient $0.01$
Value Loss Coefficient $0.5$
# PPO Epochs $8$
# PPO Minibatches $4$
# PPO Steps $400$
PPO Value Loss Clipped
PPO Value Loss Clip Value $0.2$
Reward Shaping Yes (linearly decreased over training)
Table 5: Values used for a grid search over hyperparameters governing the learning process. Finally used values appear in bold.
Description Value
Learning Rate $\eta$ [$1\cdot 10^{-4}$, $\mathbf{3\cdot 10^{-4}}$, $5\cdot 10^{-4}$, $1\cdot 10^{-3}$]
Entropy Coefficient [$\mathbf{0.01}$ $0.1$]
# PPO Steps [$256$, $\mathbf{400}$]
# Hidden Layers [$2$, $\mathbf{3}$, $4$]
Reward Shaping Annealing Steps [00, $2500000$, $5000000$, until end]
Table 6: DR hyperparameters.
Description Value
$n$ walls to place Sampled between 0015151515
$n$ onion piles to place Sampled uniformly between 11112222
$n$ plate piles to place Sampled uniformly between 11112222
$n$ pots to place Sampled uniformly between 11112222
$n$ goals to place Sampled uniformly between 11112222
Table 7: PLR specific hyperparameters in addition to the DR hyperparameters.
Description Value
UED Score MaxMC (Jiang et al., 2021a)
PLR replay probability $\rho$ $0.5$
PLR buffer size $4,000$
PLR staleness coefficient $0.3$
PLR temperature $0.1$
PLR score ranks Yes
PLR minimum fill ratio $0.5$
PLR Yes
PLR Yes
PLR force unique level Yes
Table 8: ACCEL hyperparameters in addition to the DR hyperparameters.
Description Value
UED Score MaxMC (Jiang et al., 2021a)
PLR replay probability $\rho$ $0.8$
PLR buffer size $4,000$
PLR staleness coefficient $0.3$
PLR temperature $0.1$
PLR score ranks Yes
PLR minimum fill ratio $0.5$
PLR Yes
PLR Yes
PLR force unique level Yes
ACCEL Mutation Overcooked Mutator
ACCEL $n$ mutations 20
ACCEL subsample size 4
Table 9: PAIRED hyperparameters. All PPO hyperparameters are the same between the student and the teacher. The minimax implementation follows to original one in (Dennis et al., 2020) and we stick to it too.
Description Value
$n$ walls to place Sampled between 0015151515
$n$ students 2
UED Score Relative regret (Dennis et al., 2020)
UED first wall sets budget Yes
UED noise dim $50$
PAIRED Creator OvercookedUED
We overview all hyperparameters for training in Table 4 and provide details on the hyperparameter search used in Table 5. This search was conducted on smaller single layout runs to determine reasonable values as complete runs would have been computationally infeasible. Furthermore we show the hyperparameters for each DCD method separately: DR hyperparameters in Table 6, PLR hyperparameters in Table 7, ACCEL hyperparameters in Table 8, and PAIRED hyperparameters in Table 9. DR hyperparameters govern how Overcooked levels are generated randomly and apply to all other processes in which a random level is sampled, for instance, in PLR, in which case the same hyperparameters apply.
In the experiment displayed in Figure 5 we show how policies behave when trained on all versus on only one layout and then decide on the SoftMoE architecture for our agents. Since these are considerably easier problems we train these models on fewer environment steps in total. We train the overfitting baseline for roughly $1/30$ ($12,800,00$ steps in the environment) of the experience of all UED methods and only use $1,000$ outer loops. For the Oracle baseline, we use $1/2$ of the experience ($200,000,000$ steps in the environment) and only $5,000$ outer loops. Both are trained until convergence and to speed up training we deploy 100 environment simulators. Recall that the UED methods use $30,000$ training loops and $400,000,000$ steps in the environment. Notably, in the case of population PAIRED these steps apply per student and do not include any additional steps taken in the teacher environment.
A.5 Neural network architectures
This work employs an actor-critic architecture using a separate actor and critic in which the critic is centralised for training via MAPPO (Yu et al., 2022). For the actor, the observations are of shape $h\times w\times 26$, while for the centralised critic, we concatenate the observations along the last axis to form a centralised observation, i.e. the centralised observation has shape $h\times w\times 52$ following prior work (Yu et al., 2023).
Figure 11: Default architecture featuring a convolutional encoder and an RNN.
All our networks feature a convolutional encoder $f_{c}$. This encoder always features three 2D convolutions of $32$, $64$ and $32$ channels with kernel size $3\times 3$ each and pads the input with zeros. Our default activation function is ReLU (Fukushima, 1975; Nair & Hinton, 2010) which we apply after every convolutional block. We feed the output of $f_{c}$ to a feed-forward neural network $f_{e}$ with three layers with $64$ neurons, ReLU and LayerNorm (Ba et al., 2016) applied each. $f_{e}$ takes the flattened representation produced by $f_{c}$ and produces an embedding $e\in\mathbb{R}^{b\times t\times 64}$ that we feed into a recurrent neural network (either LSTM (Hochreiter & Schmidhuber, 1997) or S5 (Smith et al., 2023)) to aggregate information along the temporal axis. We use this resulting embedding $e_{t}\in\mathbb{R}^{b\times 64}$ to produce action logits $l\in\mathbb{R}^{b\times 6}$ to parameterise a categorical distribution in the actor-network or directly produce a value $v\in\mathbb{R}^{b\times 1}$ in the critic network using a final projection layer. This architecture is inspired by previous work on Overcooked-AI, specifically (Yu et al., 2023), see Figure 11 for an overview. We also test the use of a S5 layer (Smith et al., 2023) in which case we use $2$ S5 blocks, $2$ S5 layers, use LayerNorm before the SSM block and the activation function described in the original work, i.e. $a(x)=\textrm{GELU}(x)\odot\sigma(W*\textrm{GELU}(x))$. In the case of the SoftMoE architecture, we follow the same approach as in (Obando-Ceron et al., 2024) and replace the penultimate layer with a SoftMoE layer. As in their work we use the PerConv tokenisation technique, i.e. given input $x\in\mathbb{N}^{h\times w\times 26}$ we take the output $y\in\mathbb{R}^{h\times w\times 32}$ of $f_{c}$ and construct $h\times w$ tokens with dimension $d=32$ that we then feed into the SoftMoE layer. We always use $32$ slots and $4$ experts for this layer, see (Obando-Ceron et al., 2024) for details on this layer. The resulting embedding is then passed into the two remaining linear layers before being also passed to RNN and used to produce an action or value, equivalent to the description above, also compare Figure 4.
Table 10: Number of trainable parameters in each model.
CNN-LSTM SoftMoE-LSTM CNN-S5
Parameter Count 197,254 316,102 193,670
Lastly, we describe our networks in terms of parameter count in Table 10.
A.6 Additional analysis
A.6.1 Implementation Details
Table 11: Average steps-per-second for different numbers of parallel environments measured by taking 1,000 steps with randomly sampled actions to show how our adapted Overcooked environment performs compared to a simpler single-agent UED environment.
# Parallel Envs 1 32 256 1024 4096 16384
AMaze $264$ $8,141$ $67,282$ $264,142$ $1,058,306$ $3,321,678$
OvercookedUED $151$ $4,921$ $40,011$ $156,696$ $631,836$ $2,017,526$
The OGC is implemented in Jax (Bradbury et al., 2018) and integrated into minimax (Jiang et al., 2023). As such, it can be tested with all available DCD algorithms present in minimax. To achieve this we extend minimax with runners, replay buffers etc. that are compatible with multiple agents. Building on an established library eliminates sources of error and presents users of the challenge with a familiar experience. We present the steps-per-seconds (SPS) on our setup given varying degrees of parallelism in Table 11 and compare it to the GPU-accelerated maze environment minimax includes AMaze. Given sufficiently large numbers of parallel environments, OGC can be run at hundreds of thousands of SPS. While less than AMaze, the OGC is a more fully-featured environment in which multiple agents take steps and interact.
A.6.2 Performance across levels
Table 12: Mean episode solved rate for the different methods averaged over the respective testing layouts. The best result is shown in bold. We report aggregate statistics over three random seeds. As a baseline we include an Oracle version for all architectures, which was trained on the five testing layouts directly.
Method CNN-LSTM SoftMoE-LSTM CNN-S5
DR $0.02\pm 0.0\%$ $6.31\pm 10.1\%$ $0.00\pm 0.0\%$
PLR⟂,∥ $0.00\pm 0.0\%$ $0.33\pm 0.3\%$ $0.00\pm 0.0\%$
Pop. PAIRED $0.13\pm 0.2\%$ $\mathbf{11.46\pm 2.1\%}$ $0.00\pm 0.0\%$
ACCEL $0.00\pm 0.0\%$ $0.00\pm 0.0\%$ $0.00\pm 0.0\%$
Oracle $95.40\pm 7.5\%$ $99.67\pm 0.6\%$ $97.53\pm 4.1\%$
To accompany the overall performance measured by reward in the main paper in Table 1 we also measure the mean solved rate on display it in Table 12.
A.6.3 Performance on individual levels
Table 13: Performance on all evaluation layouts. We show the mean episode reward R and the mean episode solved rate SR. The overall best result per layout is presented in bold excluding oracle results.
Layout Method CNN-LSTM SoftMoE-LSTM CNN-S5
R SR R SR R SR
Cramped DR 1.70 0.0% 1.54 0.2% 0.00 0.0%
PLR⟂,∥ 1.12 0.0% 5.02 2.1% 0.14 0.0%
Pop. PAIRED 0.8 0.0% 15.33 17.0% 0.00 0.0%
ACCEL 1.93 0.00% 2.53 0.0% 0.46 0.0%
Oracle 241.27 96.7% 245.54 100.0% 189.47 99.7%
Coord DR 0.00 0.0% 0.00 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.00 0.0% 0.00 0.0%
Pop. PAIRED 0.00 0.0% 3.27 0.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.00 0.0% 0.00 0.0%
Oracle 197.8 100.0% 204.53 100.0% 119.33 99.0%
Forced DR 0.00 0.0% 0.02 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.02 0.0% 0.02 0.0%
Pop. PAIRED 0.00 0.0% 0.00 0.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.67 0.0 % 0 0.0 %
Oracle 196.8 100.0% 204.53 100.0% 133.47 94.7%
Asymm DR 0.58 0.1% 8.64 4.4% 0.00 0.0%
PLR⟂,∥ 0.08 0.0% 0.10 0.0% 0.08 0.0%
Pop. PAIRED 2.4 0.6% 28.67 40.4% 0.00 0.0%
ACCEL 0.67 0.0% 1.00 0.0 % 0.00 0.0 %
Oracle 220.4 100.0% 277.8 98.4% 247.87 99.7%
Counter DR 0.00 0.0% 0.00 0.0% 0.00 0.0%
PLR⟂,∥ 0.00 0.0% 0.00 0.0% 0.00 0.0%
Pop. PAIRED 0.00 0.0% 0.14 0.0% 0.00 0.0%
ACCEL 0.00 0.0% 0.00 0.0 % 0.00 0.0 %
Oracle 91.2 77.3% 152.73 100.0% 84.93 94.7%
We list the performance of every individual method on every single layout in Table 13. Most notable is that some layouts are harder to learn than others. Our agents especially seem to struggle with layouts requiring more complex forms of interaction, i.e. Coordination Ring, Counter Circuit and Forced Coordination. Forced Coordination especially seems difficult to solve as no run achieves noticeable performance on it. This might be due to the specific features of the layout, i.e. agents have access to several objects and need to hand them over the counter to produce any result.
A.6.4 Population training details
(a) Coordination Ring (6x9)
(b) Cramped Room (6x9)
(c) Forced Coordination (6x9)
(d) Asymmetric Advantages (6x9)
(e) Counter Circuit (6x9)
Figure 12: Runs used for the FCP evaluation populations with random seeds 18 for the OGC with bands reporting standard error $\sigma/\sqrt{n}$. Layouts were padded to a total size of 6 x 9 to be compatible with the policies trained via DCD.
Both populations were trained over 8 random seeds. As architecture, we used a simple CNN encoder without RNN as in prior work (Zhao et al., 2023; Yu et al., 2023). To give an intuition into the performance of the members of the population, we present the training curves over all 8 seeds of training an FCP population in Figure 12. MEP was trained with the same architecture, with the same amount of experience per agent and achieved practically identical results. As in prior work (Zhao et al., 2023) we set the population entropy coefficient during training to $\alpha=0.01$.
A.6.5 Detailed results with populations
Figure 13: Ad-hoc teamwork results of the SoftMoE-LSTM policy paired with an FCP and MEP population trained on the respective layout. Error bars show standard error.
Table 14: Zero-shot results using SoftMoE-LSTM policies playing with an FCP and MEP population of experts trained on the respective layout exclusively. We report the mean episode reward and standard deviation. The best result per layout is put in bold.
Method Asymm Counter Cramped Forced Coord
FCP
Random $7.43\pm 12.19$ $8.89\pm 4.65$ $66.02\pm 38.28$ $1.95\pm 1.92$ $20.49\pm 7.82$
Stay $5.32\pm 12.07$ $0.38\pm 1.11$ $20.67\pm 33.05$ $00.00\pm 0$ 0.00 $0.95\pm 2.73$
Oracle $126.44\pm 27.13$ $22.63\pm 7.82$ $120.9\pm 10.86$ $22.08\pm 12.89$ $59.64\pm 22.17$
DR $18.18\pm 1.69$ $6.86\pm 5.27$ $65.05\pm 5.15$ $1.09\pm 0.21$ $17.88\pm 10.27$
PLR⟂,∥ $7.64\pm 0.89$ $5.60\pm 1.29$ $60.35\pm 6.89$ $1.76\pm 0.86$ $21.90\pm 1.26$
Pop. PAIRED $\mathbf{42.28\pm 19.59}$ $\mathbf{10.12\pm 1.67}$ $\mathbf{63.41\pm 9.13}$ $\mathbf{2.57\pm 1.46}$ $\mathbf{21.97\pm 2.73}$
ACCEL $8.19\pm 1.08$ $9.39\pm 3.21$ $61.67\pm 2.79$ $2.04\pm 2.37$ $17.94\pm 2.29$
MEP
Random $8.0\pm 9.12$ $22.46\pm 13.34$ $58.33\pm 34.83$ $2.55\pm 2.76$ $31.85\pm 19.69$
Stay $4.86\pm 7.21$ $5.2\pm 10.85$ $31.55\pm 47.13$ $0.0\pm 0$ 0.0 $1.53\pm 3.61$
Oracle $135.07\pm 30.27$ $39.33\pm 13.53$ $138.07\pm 10.0$ $56.1\pm 25.41$ $67.86\pm 10.89$
DR $19.32\pm 0.39$ $18.04\pm 5.75$ $\mathbf{62.77\pm 7.22}$ $1.69\pm 0.67$ $30.35\pm 4.42$
PLR⟂,∥ $7.53\pm 0.92$ $\mathbf{21.23\pm 1.91}$ $57.2\pm 4.4$ $2.45\pm 1.23$ $2.45\pm 1.23$
Pop. PAIRED $\mathbf{42.66\pm 20.31}$ $18.34\pm 4.85$ $61.64\pm 7.63$ $\mathbf{3.58\pm 1.69}$ $\mathbf{31.24\pm 5.52}$
ACCEL $9.08\pm 1.11$ $18.43\pm 1.77$ $53.02\pm 5.53$ $2.88\pm 3.31$ $28.81\pm 2.01$
Table 15: Zero-shot results using SoftMoE-LSTM policies playing with an FCP and MEP population of experts trained on the respective layout exclusively. We report the mean solved rate and standard deviation. The best result per layout is put in bold.
Method Asymm Counter Cramped Forced Coord
FCP
Random $8.52\pm 17.52\%$ $5.00\pm 6.70\%$ $69.43\pm 38.45\%$ $0.00\pm 0.00\%$ $30.89\pm 3.83\%$
Stay $6.81\pm 18.04\%$ $0.02\pm 0.14\%$ $21.75\pm 33.71\%$ $0.00\pm 0.00\%$ $0.14\pm 0.74\%$
Oracle $69.67\pm 16.39\%$ $27.39\pm 19.02\%$ $31.30\pm 20.97\%$ $92.02\pm 1.19\%$ $96.96\pm 2.23\%$
DR $24.19\pm 4.60\%$ $4.56\pm 5.32\%$ $72.11\pm 6.29\%$ $0.01\pm 0.01\%$ $23.76\pm 18.85\%$
PLR⟂,∥ $8.84\pm 1.31\%$ $2.04\pm 0.95\%$ $68.14\pm 1.21\%$ $0.11\pm 0.12\%$ $\mathbf{30.89\pm 3.83\%}$
Pop. PAIRED $\mathbf{56.91\pm 25.08\%}$ $6.07\pm 2.54\%$ $\mathbf{72.48\pm 6.14\%}$ $0.2\pm 0.41\%$ $30.16\pm 7.7\%$
ACCEL $8.79\pm 1.59\%$ $\mathbf{6.45\pm 5.18\%}$ $68.09\pm 2.16\%$ $\mathbf{0.51\pm 0.89\%}$ $20.67\pm 5.18\%$
MEP
Random $9.25\pm 2.02\%$ $36.04\pm 4.38\%$ $67.75\pm 5.48\%$ $0.00\pm 0.00\%$ $54.9\pm 5.55\%$
Stay $4.91\pm 1.46\%$ $5.85\pm 2.71\%$ $29.56\pm 5.92\%$ $0.00\pm 0.00\%$ $1.02\pm 0.51\%$
Oracle $91.02\pm 1.12\%$ $52.60\pm 11.37\%$ $96.86\pm 2.27\%$ $56.16\pm 21.85\%$ $75.23\pm 0.91\%$
DR $26.34\pm 3.55\%$ $27.41\pm 10.31\%$ $70.78\pm 4.23\%$ $0.05\pm 0.07\%$ $50.07\pm 6.67\%$
PLR⟂,∥ $8.24\pm 1.28\%$ $\mathbf{33.76\pm 4.89\%}$ $65.38\pm 4.55\%$ $0.28\pm 0.41\%$ $\mathbf{50.97\pm 4.01\%}$
Pop. PAIRED $\mathbf{57.43\pm 26.49\%}$ $24.72\pm 10.42\%$ $\mathbf{72.97\pm 7.6\%}$ $0.4\pm 0.5\%$ $50.64\pm 7.5\%$
ACCEL $9.16\pm 2.14\%$ $25.91\pm 4.56\%$ $64.31\pm 4.01\%$ $\mathbf{1.38\pm 2.33\%}$ $49.23\pm 2.89\%$
We present detailed zero-shot cooperation results per layout in Table 14 and 15. As indicated through the averaged performance discussed in the main text, we also find that PAIRED performs best on four of the five individual layouts in terms of zero-shot cooperation.
A.7 Error Analysis
Figure 14: An illustration of the circular evaluation levels; we move the kitchen around the sides and vary the size.
We hypothesise that agents may fail to generalise since observations are hard to generalise from. To test this we design testing layouts that rotate and mirror features of the environment to look for systemic failures, i.e. cases in which an agent does well in one environment but not its mirrored version. While methods generally differed in how well they performed along the same line as previous analysis suggests, the low standard deviations show that any given method performs similarly well on each of the 24 layouts, see Table 2.
A.7.1 Training curves and evaluation
Figure 15: Returns in training and evaluation levels over the duration of training for our SoftMoE architecture.
Figure 16: Returns in training and evaluation levels over the duration of training for our S5 architecture.
Figure 17: Returns in training and evaluation levels over the duration of training for our CNN-LSTM architecture.
In Figure 15, Figure 16 and Figure 17 we show the returns of our agent during training in seen training levels, as well as the five unseen evaluation levels. The results for the SoftMoE architecture are displayed in Figure 15, the results for the S5 in Figure 16 and the results for the CNN-LSTM in Figure 17. Interestingly, while (SoftMoE) PAIRED performs the best in our evaluations it does not reach the highest training returns, instead it achieves the highest training return while keeping the generalisation gap small.